What is the common point between lines y = ½ x + 4 and y = x – 4?
1
Expert's answer
2013-07-12T07:46:30-0400
Common point belongs to both given lines. To find it we should equate given expressions for y:
1/2 x+4=x-4
1/2x=8
x=16
Now we need to find corresponding y. We can use either of the given equations for that:
y(16)=16-4=12
Thus, we found common point of the given lines, it's (16,12)